# SPECIAL RELATIVITY

### v9

@@ 13 Relativistic Particle Dynamics
In previous chapters we considered the kinematics of the theory of relativity, which concerns the study of the four vectors of positions, velocity and acceleration. Kinematics studies trajectories as geometrical objects, independently of their causes. This means that it is not possible to predict the trajectory of a particle evolving under a given dynamical field using just a kinematic treatment. In dynamics we consider the effect of interaction on motion.
@@ 13.1 Manifestly Covariant Particle Dynamics
Dynamics equations can be expressed as tensor equations in Minkowski space-time. When coordinates are chosen, one may work with components, instead of geometric objects. Relying on the geometric structure of Minkowski space-time, one can define the class of inertial frames and can adopt a Lorentz frame with orthonormal basis vectors for any given inertial frame. In any Lorentz coordinate system the law of motion becomes
@@
where here the particle’s mass and charge are denoted by $m$ and $e$ respectively. The electromagnetic field is described by a second-rank, antisymmetric tensor with components $F^{\mu\nu}$. The coordinate-independent proper time $\tau$ is a parameter describing the evolution of physical system under the relativistic laws of motion, Eq. (20).
The covariant equation of motion for a relativistic charged particle under the action of the four-force $K_{\mu}=eF^{\mu\nu}dx_{\nu}/d\tau$ in the Lorentz lab frame, Eq.(20), is a relativistic ”generalization” of the Newton’s second law. The three-dimensional Newton second law $md\vec{v}/dt=\vec{f}$ can always be used in the instantaneous Lorentz comoving frame. Relativistic ”generalization” means that the previous three independent equations expressing Newton second law are be embedded into the four-dimensional Minkowski space-time (1).
@@
Having written down the motion equation in a 4-vector form, Eq.(20), and determined the components of the 4-force, we satisfied the principle of relativity for one thing, and, for another, we obtained the four components of the equation of particle motion. This is covariant relativistic generalization of the three dimensional Newton’s equation of motion which is based on particle proper time as the evolution parameter.
We next wish to describe a particle motion in the Lorentz lab frame using the lab time $t$ as evolution parameter. First, we examine the reasoning presented in textbooks. Let us determine the first three spatial components of the 4-force. We consider for this the spatial part of the dynamics equation, Eq.(20): $\vec{Q}=(dt/d\tau)d(m\gamma\vec{v})/dt=\gamma d(m\gamma\vec{v})/dt$. The prefactor $\gamma$ arises from the change of the evolution variable from the proper time $\tau$, which is natural since $\vec{Q}$ is the space part of a four-vector, to the lab frame time $t$, which is needed to introduce the usual force three-vector $\vec{f}$: $\vec{Q}=\gamma\vec{f}$. Written explicitly, the relativistic form of the three-force is
@@
According to the non-covariant (3+1) approach we seek for the initial value solution to these equations. Using explicit expression for Lorentz force we find that the three equations Eq.(21) automatically imply the constraint Eq.(23), once this is satisfied initially at $t=0$. In the (3+1) approach, the four equations of motion ”split up” into (3+1) equations and we have no mixture of space and time parts of the dynamics equation Eq.(20). This approach to relativistic particle dynamics relies on the use of three independent equations of motion Eq.(21) for three independent coordinates and velocities, ”independent” meaning that equation Eq.(22) (and constraint Eq.(23)) are automatically satisfied.
One could expect that the particle’s trajectory in the lab frame, following from the previous reasoning $\vec{x}(t)$, should be identified with $\vec{x}_{cov}(t)$. However, paradoxical result are obtained by doing so. In particular, the trajectory $\vec{x}(t)$ does not include relativistic kinematics effects. In the non-covariant (3+1) approach, the solution of the dynamics problem in the lab frame makes no reference to Lorentz transformations. This means that, for instance, within the lab frame the motion of particles in constant magnetic field looks precisely the same as predicted by Newtonian kinematics: relativistic effects do not have a place in this description. In conventional particle tracking a particle trajectory $\vec{x}(t)$ can be seen from the lab frame as the result of successive Galileo boosts that track the motion of the accelerated (in a constant magnetic field) particle. The usual Galileo rule for addition of velocities is used to determine the Galileo boosts tracking a particular particle, instant after instant, along its motion along the curved trajectory.
The old kinematics is especially surprising, because we are based on the use of the covariant approach. Where does it comes from? The previous commonly accepted derivation of the equations for the particle motion in the three dimensional space from the covariant equation Eq.(20) includes one delicate point. In Eq.(21) and Eq.(22) the restriction $\vec{p}=m\vec{v}/\sqrt{1-v^{2}/c^{2}}$ has already been imposed. One might well wonder why, because in the accepted covariant approach, the solution of the dynamics problem for the momentum in the lab frame makes no reference to the three-dimensional velocity. In fact, equation Eq.(20) tells us that the force is the rate of change of the momentum $\vec{p}$, but does not tell us how momentum varies with speed. The four-velocity cannot be decomposed into $u=(c\gamma,\vec{v}\gamma)$ when we deal with a particle accelerating along a curved trajectory in the Lorentz lab frame.
@@
It is important to stress in this point that the situation when only one clock in the comoving frame is involved in dynamics cannot be realized. The Newton law can be written down in the proper frame only when a space-time coordinate system $(x^{\prime},y^{\prime},z^{\prime},\tau)$ has been specified. The type of clock synchronization which provides the time coordinate in the Newton equation has never been discussed in textbooks. It is clear that without an answer to the question about the method of distant clock synchronization used, not only the concept of acceleration, but also the dynamics law has no physical meaning. A proper frame approach to relativistic particle dynamics is forcefully based on a definite (Einstein) coordinatization assumption. After this, the dynamics theory states that the equation of motion in the proper frame is $md^{2}\vec{x^{\prime}}/d\tau^{2}=\vec{f}^{\prime}$. It should be stressed that in the case of velocity increment $\Delta\vec{v^{\prime}}=\Delta\vec{x^{\prime}}/\Delta\tau$ we also deal with the distant events. It can be said with some abuse of language that 3-velocity vector is always a ”spatially extended” object. Let us return to our consideration of the relation $d\tau=dt/\gamma$ between the $\tau$ and lab time $t$. The calculation carried out in the case of an spatially extended object shows that the temporal coincidence of two distant events has absolute character: $\Delta\tau=0$ implies $\Delta t=0$.
@@
It should be note that usually the notation ”$\tau$” arose from the proper time. Now let us recall the standard concept of the object’s proper time. Let a point-like object move uniformly and rectilinearly relative to the lab frame $K$. The proper frame $K^{\prime}$ can be fixed to the moving object. The object is at rest in this frame, so that events happening with this object are registered by one clock. This clock counts the proper time at the point where the object is located. As we already know, the proper time on moving object flows slower than the time $t$. This phenomena was called time dilation. The proper time can also be introduced for a point particle moving with acceleration. This standard interpretation of the proper time $\tau$ has evidently nothing to do with dynamics and one may wonder where this contradiction existing between the name and the content of time in dynamic law come from.
@@
In order to avoid being to abstract for to long we have given one example. Let us discuss the case in which the velocity increment is not in the direction of the uniform motion. For example, there may be a particle which is just moving ”upward” with velocity increment $\Delta y^{\prime}/\Delta\tau$ and the lab system is moving ”horizontally”. Now, according to textbooks, we simply using relation $d\tau=dt/\gamma$ with the standard result $\Delta v_{y}=\Delta v^{\prime}_{y}/\gamma$. However, paradoxical result are obtained by doing so. In particular, the trajectory of the particle does not include mixture of positions and time. One of the important conclusions of the discussion presented above in the Chapter 7 is that there is no absolute notion of simultaneity in the theory of special relativity. Therefore there is no notion of an instantaneous 3-space. A coordinate axes of reference frame $K^{\prime}$ rotating about the axes of a Lorentz lab frame. The standard presentation of the velocity increment transformation as the relation $\Delta v_{y}=\Delta v^{\prime}_{y}/\gamma$ is based on the hidden assumption that $y^{\prime}$-axis and $y$-axis are parallel. On the other hand Lorentz transformation makes axes ($y^{\prime}$ and $y$) oblique-angled due to Wigner rotation. A difference between non-covariant and covariant point particle trajectories follows from this immediately (We will discuss the transformation of velocities in greater detail in Section 13.5).
@@ 13.3 Incorrect Expansion of the Relation $d\tau=dt/\gamma$ to an Arbitrary Motion
How can such an unusual momentum-velocity relation come about? We know that the components of momentum four-vector $p_{\mu}=(E/c,\vec{p})$ behave under transformations from one Lorentz frame to another, exactly in the same manner as the component of the four-vector event $x=(x_{0},\vec{x})$. Surprises can surely be expected when we return from the four-vectors language to the three-dimensional velocity vector $\vec{v}$, which can be represented in terms of the components of four-vector as $\vec{v}=d\vec{x}/dx_{0}$. In contrast with the pseudo-Euclidean four-velocity space, the relativistic three-velocity space is a three-dimensional space with constant negative curvature, i.e. three-dimensional space with Lobachevsky geometry.
@@
We can see why by examine the transformation of the three velocity in the theory of relativity. For a rectilinear motion, this transformation is performed as $v=(v^{\prime}+V)/(1+v^{\prime}V/c^{2})$. The relativistic factor $1/\sqrt{1-v^{2}/c^{2}}$ is given by: $1/\sqrt{1-v^{2}/c^{2}}=(1+v^{\prime}V/c^{2})/(\sqrt{1-v^{\prime 2}/c^{2}}\sqrt{1-V^{2}/c^{2}})$. The new momentum is then simply $mv$ times the above expression. But we want to express the new momentum in terms of the primed momentum and energy, and we note that $p=(p^{\prime}+\mathcal{E}^{\prime}V/c^{2})/\sqrt{1-V^{2}/c^{2}}$. Thus, for a rectilinear motion, the combination of Einstein addition law for parallel velocities and the usual momentum-velocity relation is consistent with the covariant momentum transformation. The dynamical evolution of the particle according to the non-covariant particle tracking may be considered as the result of successive infinitesimal Lorentz transformations. The corrected Newton law is valid for each step. In other words, the two observers with rectilinear trajectories have common 3-space. This result was incorrectly extended to an arbitrary trajectory.
We already know that the collinear Lorentz boosts (similar to the collinear Galilean boosts) form a commutative group. This means that the resultant of successive collinear Lorentz boosts is independent of the transformation order. On the contrary, Lorentz boosts in different directions do not commute and do not make up a group. In contrast, the set of Galilean boosts always form a group. As well-known, the composition of non-collinear Lorentz boosts is equivalent to a Lorentz boost followed by a spatial rotation. This Wigner rotation is relativistic effect that does not have a non-covariant analogue. One of the consequences of non-commutativity of non-collinear Lorentz boosts is the unusual momentum-velocity relation $\vec{p}_{cov}\neq m\vec{v}_{cov}/\sqrt{1-v^{2}_{cov}/c^{2}}$, which also does not have any non-covariant analogue.
The theory of relativity shows that the unusual momentum-velocity relation discussed above is related with the acceleration along curved trajectories. In this case there is no notion of a common ordinary space and there is a difference between covariant and non-covariant particle trajectories. Only the solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle trajectories in the lab frame. A closer analysis of the concept of velocity, i.e. a discussion of the methods by which a time coordinate can actually be assigned in the lab frame, opens up the possibility of a description of such physical phenomena as radiation from a relativistic electron accelerating along a curved trajectory in accordance with the theory of relativity.
@@ 13.4 The Usual Integration of the 4D Covariant Equation of Motion
Attempts to solve the dynamics equation Eq.(20) in manifestly covariant form can be found in the literature. The trajectory which is found does not include relativistic kinematics effects. Therefore, it cannot be identified with $\vec{x}_{cov}(t)$ even if, at first glance, it appears to be derived following covariant prescription.
@@
First, we analyze textbooks presentation of the integration of the covariant equation of motion. Consider, for example, the motion of a particle in a given electromagnetic field. The simplest case, of great practical importance, is that of an uniform electromagnetic field meaning that $F^{\mu\nu}$ is constant on the whole space-time region of interest. In particular we consider the motion of a particle in a constant homogeneous magnetic field, specified by tensor components $F^{\mu\nu}=B(e^{\mu}_{2}e^{\nu}_{3}-e^{\nu}_{2}e^{\mu}_{3})$ where $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are orthonormal space like basis vectors $e^{2}_{2}=e^{2}_{3}=-1$, $e_{2}\cdot e_{3}=0$. In the lab frame of reference where $e^{\mu}_{0}$ is taken as the time axis, and $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are space vectors the field is indeed purely magnetic, of magnitude $B$ and parallel to the $e_{1}$ axis. Let us set the initial four-velocity $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$, where $v$ is the initial particle’s velocity relative to the lab observer along the axis $e_{2}$ at the instant $\tau=0$, and $\gamma=1/\sqrt{1-v^{2}/c^{2}}$. The components of the equation of motion are then $du^{(0)}/d\tau=0$, $du^{(1)}/d\tau=0$, $du^{(2)}/d\tau=-eBu^{(3)}/(mc)$, $du^{(3)}/d\tau=eBu^{(2)}/(mc)$. We seek for the initial value solution to these equations as done in the existing literature (2). A distinctive feature of the initial value problem in relativistic mechanics, is that the dynamics is always constrained. In fact, the evolution of the particle is subject to $mdu^{\mu}/d\tau=eF^{\mu\nu}u_{\nu}$, but also to the constraint $u^{2}=c^{2}$. However, such a condition can be weakened requiring its validity at certain values of $\tau$ only, let us say initially, at $\tau=0$. Therefore, if $Y(\tau)$ vanishes initially, i.e. $Y(0)=0$, then $Y(\tau)=0$ at any $\tau$. In other words, the differential Lorentz-force equation implies the constraint $u^{2}=c^{2}$ once this is satisfied initially. Integrating with respect to the proper time we have $u^{\mu}(\tau)=\gamma e^{\mu}_{0}+\gamma v[e^{\mu}_{2}\cos(\omega\tau)+e^{\mu}_{3}\sin(\omega\tau)]$ where $\omega=eB/(mc)$. We see that $\gamma$ is constant with time, meaning that the energy of a charged particle moving in a constant magnetic field is constant. After two successive integrations we have $X^{\mu}(\tau)=X^{\mu}(0)+\gamma c\tau e^{\mu}_{0}+R[e^{\mu}_{2}\sin(\omega\tau)-e^{\mu}_{3}\cos(\omega\tau)]$ where $R=\gamma v/\omega$. This enables us to find the time dependence $[0,X^{(2)}(t),X^{(3)}(t)]$ of the particle’s position since $t/\gamma=\tau$. From this solution of the equation of motion we conclude that the motion of a charged particle in a constant magnetic field is a uniform circular motion.
@@
One could expect that the particle’s trajectory in the lab frame, following from the previous reasoning $[0,X^{(2)}(t),X^{(3)}(t)]$, should be identified with $\vec{x}_{cov}(t)$. However, paradoxical result are obtained by doing so. In particular, the trajectory $[0,X^{(2)}(t),X^{(3)}(t)]$ does not include relativistic kinematics effects. We found that the usual integration of the four-dimensional covariant equation of motion Eq.(20) gives particle trajectory which looks precisely the same as in Newton dynamics and kinematics. The trajectory of the electron does not include relativistic effects and the Galilean vectorial law of addition of velocities is actually used.
@@
So we must have made a mistake. Notice that we are using textbooks to solve covariant equation of motion. We did not make a computational mistake in our integrations, but rather a conceptual one. We must say immediately that there is no objection to the first integration of Eq.(20) from initial conditions over proper time $\tau$. With this, we find the four-momentum. The momentum has exact objective meaning i.e. it is convention-invariant. What must be recognized is that the concept of velocity is only introduced in the second integration step. However, in accepted covariant approach, the solution of the dynamics problem for the momentum in the lab frame makes no reference to three-dimensional velocity. In fact, the initial condition which we used is $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$ and includes $\gamma c$ and $\gamma v$, which are actually notations for the time and space parts of the initial four-momentum. The three-dimensional trajectory and respectively velocity, which are convention-dependent, are only found after the second integration step. Then, where does the old kinematics comes from? The second integration was performed using the relation $d\tau=dt/\gamma$. It is only after we have made those replacement for $d\tau$ that we obtain the usual formula for conventional (non-covariant) trajectory for an electron in a constant magnetic field.

### v10

@@ 16.1 Manifestly Covariant Particle Dynamics
Dynamics equations can be expressed as tensor equations in Minkowski space-time. When coordinates are chosen, one may work with components, instead of geometric objects. Relying on the geometric structure of Minkowski space-time, one can define the class of inertial frames and can adopt a Lorentz frame with orthonormal basis vectors for any given inertial frame. In any Lorentz coordinate system, the law of motion becomes
@@
where here the particle’s mass and charge are denoted by $m$ and $e$ respectively. The electromagnetic field is described by a second-rank, antisymmetric tensor with components $F^{\mu\nu}$. The coordinate-independent proper time $\tau$ is a parameter describing the evolution of the physical system under the relativistic laws of motion, Eq. (21).
The covariant equation of motion for a relativistic charged particle under the action of the four-force $K_{\mu}=eF^{\mu\nu}dx_{\nu}/d\tau$ in the Lorentz lab frame, Eq.(21), is a relativistic ”generalization” of Newton’s second law. The three-dimensional Newton’s second law $md\vec{v}/dt=\vec{f}$ can always be used in the instantaneous Lorentz comoving frame. Relativistic ”generalization” means that the previous three independent equations expressing Newton’s second law are embedded into the four-dimensional Minkowski space-time  [^59] The idea of embedding is based on the fact that the usual Newton’s second law can always be used in any Lorentz frame where the particle, whose motion we want to describe, is at rest [63]. In other words, if an instantaneously comoving Lorentz frame is given at some instant, one can precisely predict the evolution of the particle in this frame during an infinitesimal time interval. In geometric language, Newton’s law is strict on a hyperplane perpendicular to the world line. However, the hyperplane tilts together with its normal $u_{\mu}$ as one moves along the world line. For the embedding, we need an operator $\hat{P}_{\perp}$ that continually projects vectors of Minkowski space on hyperplanes perpendicular to the world line. The desired operator is $(\hat{P}_{\perp})_{\mu\nu}=\eta_{\mu\nu}-u_{\mu}u_{\nu}/u^{2}$ [63]. The presence of the projector operator $\hat{P}_{\perp}$ suggests that we have only three independent equations..
@@
Having written down the motion equation in a 4-vector form, Eq.(21), and determined the components of the 4-force, we satisfied the principle of relativity for one thing, and, for another, we obtained the four components of the equation of particle motion. This is a covariant relativistic generalization of the three-dimensional Newton’s equation of motion based on particle proper time as the evolution parameter.
We next wish to describe a particle motion in the Lorentz lab frame using the lab time $t$ as an evolution parameter. First, we examine the reasoning presented in textbooks. Let us determine the first three spatial components of the 4-force. We consider for this the spatial part of the dynamics equation, Eq.(21): $\vec{Q}=(dt/d\tau)d(m\gamma\vec{v})/dt=\gamma d(m\gamma\vec{v})/dt$. The prefactor $\gamma$ arises from the change of the evolution variable from the proper time $\tau$, which is natural since $\vec{Q}$ is the space part of a four-vector, to the lab frame time $t$, which is needed to introduce the usual force three-vector $\vec{f}$: $\vec{Q}=\gamma\vec{f}$. Written explicitly, the relativistic form of the three-force is
@@
According to the non-covariant (3+1) approach we seek the initial value solution to these equations. Using the explicit expression for Lorentz force we find that the three equations Eq.(22) automatically imply the constraint Eq.(24), once this is satisfied initially at $t=0$. In the (3+1) approach, the four equations of motion ”split up” into (3+1) equations and we have no mixture of space and time parts of the dynamics equation Eq.(21). This approach to relativistic particle dynamics relies on the use of three independent equations of motion Eq.(22) for three independent coordinates and velocities, ”independent” meaning that equation Eq.(23) (and constraint Eq.(24)) are automatically satisfied.
One could expect that the particle’s trajectory in the lab frame, following from the previous reasoning $\vec{x}(t)$, should be identified with $\vec{x}_{cov}(t)$. However, paradoxical results are obtained by doing so. In particular, the trajectory $\vec{x}(t)$ does not include relativistic kinematics effects. In the non-covariant (3+1) approach, the solution of the dynamics problem in the lab frame makes no reference to Lorentz transformations. This means that, for instance, within the lab frame the motion of particles in a constant magnetic field looks precisely the same as predicted by Newtonian kinematics: relativistic effects do not have a place in this description. In conventional particle tracking a particle trajectory $\vec{x}(t)$ can be seen from the lab frame as the result of successive Galileo boosts that track the motion of the accelerated (in a constant magnetic field) particle. The usual Galileo rule for the addition of velocities is used to determine the Galileo boosts tracking a particular particle, instant after instant, along its motion along the curved trajectory.
The old kinematics is especially surprising, because we are based on the use of the covariant approach. Where does it come from? The previous commonly accepted derivation of the equations for the particle motion in the three-dimensional space from the covariant equation Eq.(21) includes one delicate point. In Eq.(22) and Eq.(23) the restriction $\vec{p}=m\vec{v}/\sqrt{1-v^{2}/c^{2}}$ has already been imposed. One might well wonder why because in the accepted covariant approach, the solution of the dynamics problem for the momentum in the lab frame makes no reference to the three-dimensional velocity. Equation Eq.(21) tells us that the force is the rate of change of the momentum $\vec{p}$, but does not tell us how momentum varies with speed. The four-velocity cannot be decomposed into $u=(c\gamma,\vec{v}\gamma)$ when we deal with a particle accelerating along a curved trajectory in the Lorentz lab frame.
@@
It is important to stress at this point that the situation when only one clock in the comoving frame is involved in dynamics cannot be realized. The Newton law can be written down in the proper frame only when a space-time coordinate system $(x^{\prime},y^{\prime},z^{\prime},\tau)$ has been specified. The type of clock synchronization which provides the time coordinate in the Newton equation has never been discussed in textbooks. It is clear that without an answer to the question about the method of distant clock synchronization used, not only the concept of acceleration but also the dynamics law has no physical meaning. A proper frame approach to relativistic particle dynamics is forcefully based on a definite (Einstein) coordinatization assumption. After this, the dynamics theory states that the equation of motion in the proper frame is $md^{2}\vec{x^{\prime}}/d\tau^{2}=\vec{f}^{\prime}$. It should be stressed that in the case of velocity increment $\Delta\vec{v^{\prime}}=\Delta\vec{x^{\prime}}/\Delta\tau$, we also deal with the distant events. It can be said with some abuse of language that a 3-velocity vector is always a ”spatially extended” object. Let us return to our consideration of the relation $d\tau=dt/\gamma$ between the $\tau$ and lab time $t$. The calculation carried out in the case of a spatially extended object shows that the temporal coincidence of two distant events has absolute character: $\Delta\tau=0$ implies $\Delta t=0$.
@@
It should be noted that usually the notation ”$\tau$” arose from the proper time. Now let us recall the standard concept of the object’s proper time. Let a point-like object move uniformly and rectilinearly relative to the lab frame $K$. The proper frame $K^{\prime}$ can be fixed to the moving object. The object is at rest in this frame, so that events happening with this object are registered by one clock. This clock counts the proper time at the point where the object is located. As we already know, the proper time on moving object flows slower than the time $t$. This phenomenon was called time dilation. The proper time can also be introduced for a point particle moving with acceleration. This standard interpretation of the proper time $\tau$ has nothing to do with dynamics and one may wonder where this contradiction existing between the name and the content of time in dynamic law comes from.
@@
To avoid being to abstract for too long we have given one example. Let us discuss the case where the velocity increment is not in the direction of the uniform motion. For example, there may be a particle which is moving ”upward” with velocity increment $\Delta y^{\prime}/\Delta\tau$ in the frame $S_{n}$ and the $S_{n}$ system is moving ”horizontally”. Now, according to textbooks, we simply use relation $d\tau=dt/\gamma$ with the standard result $\Delta v_{y}=\Delta v^{\prime}_{y}/\gamma$. However, paradoxical results are obtained by doing so. In particular, the trajectory of the particle does not include a mixture of positions and time. One of the important conclusions of the discussion presented above in Section 6 is that there is no absolute notion of simultaneity in the theory of special relativity. Therefore there is no notion of an instantaneous 3-space. A coordinate axes of reference frame $K^{\prime}$ rotating about the axes of a Lorentz lab frame. The standard presentation of the velocity increment transformation as the relation $\Delta v_{y}=\Delta v^{\prime}_{y}/\gamma$ is based on the hidden assumption that $y^{\prime}$-axis and $y$-axis are parallel. On the other hand, Lorentz transformation makes axes ($y^{\prime}$ and $y$) oblique-angled due to Wigner rotation. A difference between non-covariant and covariant point particle trajectories follows from this immediately (We will discuss the transformation of velocities in greater detail in Section 16.5).
@@ 16.3 Incorrect Expansion of the Relation $d\tau=dt/\gamma$ to an Arbitrary Motion
How can such an unusual momentum-velocity relation come about? We know that the components of momentum four-vector $p_{\mu}=(E/c,\vec{p})$ behave under transformations from one Lorentz frame to another, exactly in the same manner as the component of the four-vector event $x=(x_{0},\vec{x})$. Surprises can surely be expected when we return from the four-vectors language to the three-dimensional velocity vector $\vec{v}$, which can be represented in terms of the components of four-vector as $\vec{v}=d\vec{x}/dx_{0}$.
@@
We can see why by examining the transformation of the 3-velocity in the theory of relativity. For a rectilinear motion, this transformation is performed as $v=(v^{\prime}+V)/(1+v^{\prime}V/c^{2})$. The relativistic factor $1/\sqrt{1-v^{2}/c^{2}}$ is given by: $1/\sqrt{1-v^{2}/c^{2}}=(1+v^{\prime}V/c^{2})/(\sqrt{1-v^{\prime 2}/c^{2}}\sqrt{1-V^{2}/c^{2}})$. The new momentum is then simply $mv$ times the above expression. But we want to express the new momentum in terms of the primed momentum and energy, and we note that $p=(p^{\prime}+\mathcal{E}^{\prime}V/c^{2})/\sqrt{1-V^{2}/c^{2}}$. Thus, for a rectilinear motion, the combination of Einstein’s addition law for parallel velocities and the usual momentum-velocity relation is consistent with the covariant momentum transformation.
We already know that the collinear Lorentz boosts (similar to the collinear Galilean boosts) form a commutative group. This means that the resultant of successive collinear Lorentz boosts is independent of the transformation order. On the contrary, Lorentz boosts in different directions do not commute and do not make up a group. In contrast, the set of Galilean boosts always form a group.
As well-known, the composition of non-collinear Lorentz boosts is equivalent to a Lorentz boost followed by a spatial rotation. This Wigner rotation is a relativistic effect that does not have a non-covariant analogue. One of the consequences of non-commutativity of non-collinear Lorentz boosts is the unusual momentum-velocity relation $\vec{p}_{cov}\neq m\vec{v}_{cov}/\sqrt{1-v^{2}_{cov}/c^{2}}$, which also does not have any non-covariant analog. The theory of relativity shows that the unusual momentum-velocity relation discussed above is related to the acceleration along curved trajectories. In this case, there is no notion of a common ordinary space and there is a difference between covariant and non-covariant particle trajectories.
@@ 16.4 The Usual Integration of the 4D Covariant Equation of Motion
Attempts to solve the dynamics equation Eq.(21) in manifestly covariant form can be found in the literature. The trajectory that is found does not include relativistic kinematics effects. Therefore, it cannot be identified with $\vec{x}_{cov}(t)$ even if, at first glance, it appears to be derived following covariant prescription.
@@
First, we analyze the textbook’s presentation of the integration of the covariant equation of motion. Consider, for example, the motion of a particle in a given electromagnetic field. The simplest case, of great practical importance, is that of a uniform electromagnetic field meaning that $F^{\mu\nu}$ is constant on the whole space-time region of interest. In particular, we consider the motion of a particle in a constant homogeneous magnetic field, specified by tensor components $F^{\mu\nu}=B(e^{\mu}_{2}e^{\nu}_{3}-e^{\nu}_{2}e^{\mu}_{3})$ where $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are orthonormal space like basis vectors $e^{2}_{2}=e^{2}_{3}=-1$, $e_{2}\cdot e_{3}=0$. In the lab frame of reference where $e^{\mu}_{0}$ is taken as the time axis, and $e^{\mu}_{2}$ and $e^{\mu}_{3}$ are space vectors the field is indeed purely magnetic, of magnitude $B$ and parallel to the $e_{1}$ axis. Let us set the initial four-velocity $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$, where $v$ is the initial particle’s velocity relative to the lab observer along the axis $e_{2}$ at the instant $\tau=0$, and $\gamma=1/\sqrt{1-v^{2}/c^{2}}$. The components of the equation of motion are then $du^{(0)}/d\tau=0$, $du^{(1)}/d\tau=0$, $du^{(2)}/d\tau=-eBu^{(3)}/(mc)$, $du^{(3)}/d\tau=eBu^{(2)}/(mc)$. We seek the initial value solution to these equations as done in the existing literature 606060 From the solution of the covariant equation of motion authors of textbooks [5, 63, 64] conclude that the covariant motion of a charged particle in a constant magnetic field is a uniform circular motion. The trajectory of the particle does not include relativistic kinematics effects and the Galilean vectorial law of addition of velocities is actually used. The old kinematics comes from the relation $d\tau=dt/\gamma$. It is only after the authors of textbooks have make those replacement for $d\tau$ that authors obtain the usual formula for non-covariant trajectory.. A distinctive feature of the initial value problem in relativistic mechanics is that the dynamics is always constrained. The evolution of the particle is subject to $mdu^{\mu}/d\tau=eF^{\mu\nu}u_{\nu}$, but also to the constraint $u^{2}=c^{2}$. However, such a condition can be weakened requiring its validity at certain values of $\tau$ only, let us say initially, at $\tau=0$. Therefore, if $Y(\tau)=u^{2}-c^{2}$ vanishes initially, i.e. $Y(0)=0$, then $Y(\tau)=0$ at any $\tau$. In other words, the differential Lorentz-force equation implies the constraint $u^{2}=c^{2}$ once this is satisfied initially. Integrating with respect to the proper time we have $u^{\mu}(\tau)=\gamma e^{\mu}_{0}+\gamma v[e^{\mu}_{2}\cos(\omega\tau)+e^{\mu}_{3}\sin(\omega\tau)]$ where $\omega=eB/(mc)$. We see that $\gamma$ is constant with time, meaning that the energy of a charged particle moving in a constant magnetic field is constant. After two successive integrations we have $X^{\mu}(\tau)=X^{\mu}(0)+\gamma c\tau e^{\mu}_{0}+R[e^{\mu}_{2}\sin(\omega\tau)-e^{\mu}_{3}\cos(\omega\tau)]$ where $R=\gamma v/\omega$. This enables us to find the time dependence $[0,X^{(2)}(t),X^{(3)}(t)]$ of the particle’s position since $t/\gamma=\tau$. From this solution of the equation of motion we conclude that the motion of a charged particle in a constant magnetic field is a uniform circular motion.
@@
One could expect that the particle’s trajectory in the lab frame, following from the previous reasoning $[0,X^{(2)}(t),X^{(3)}(t)]$, should be identified with $\vec{x}_{cov}(t)$. However, the trajectory $[0,X^{(2)}(t),X^{(3)}(t)]$ does not include relativistic kinematics effects. We found that the usual integration of the four-dimensional covariant equation of motion Eq.(21) gives particle trajectory which looks precisely the same as in Newton’s kinematics and dynamics. The trajectory of the electron does not include relativistic effects and the Galilean vectorial law of addition of velocities is actually used.
@@
So we must have made a mistake. Notice that we are using textbooks to solve covariant equations of motion. We did not make a computational mistake in our integrations, but rather a conceptual one. We must say immediately that there is no objection to the first integration of Eq.(21) from initial conditions over proper time $\tau$. With this, we find the four-momentum. The momentum has exact objective meaning i.e. it is convention-invariant. What must be recognized is that the concept of velocity is only introduced in the second integration step. However, in the accepted covariant approach, the solution of the dynamics problem for the momentum in the lab frame makes no reference to three-dimensional velocity. The initial condition that we used is $u^{\mu}(0)=\gamma ce^{\mu}_{0}+\gamma ve^{\mu}_{2}$ and includes $\gamma c$ and $\gamma v$, which are actually notations for the time and space parts of the initial four-momentum. The three-dimensional trajectory and velocity, which are convention-dependent, are only found after the second integration step. Then, where does the old kinematics come from? The second integration was performed using the relation $d\tau=dt/\gamma$. It is only after we have made those replacement for $d\tau$ that we obtain the usual formula for conventional (non-covariant) trajectory for an electron in a constant magnetic field.